OpenStudy (anonymous):

If the roots of the equation x^2 + px + q = 0 differ from the roots of the equation x^2+qx+p=0 by the same quantity, then (p is not equal to q) - a) p+q+1=0 c) p+q+4=0 b)p+q+2=0 d) None of these

OpenStudy (shubhamsrg):

whats the ques ?? o.O

OpenStudy (anonymous):

@shubhamsrg see on top

OpenStudy (shubhamsrg):

it says "then .." then what ?

OpenStudy (anonymous):

then you have 4 options

OpenStudy (shubhamsrg):

hmm,,what are the options supposed to be equal to ? then what? find the value of what? eh? you getting me?

OpenStudy (anonymous):

sorry all options = 0

OpenStudy (shubhamsrg):

using @apoorvk 's work, (p+q)(p-q) = -4(p-q) since p is not equal to q, we can cancel out (p-q) => p+q = -4 => b)

OpenStudy (apoorvk):

oh so sorry there were a few mistakes up there in the first one. hmm.. Let the first quadratic's roots be 'a' and 'b', and second quadratic's roots be 'c' and 'd' So, \(|a-b|=|c-d|\) For the sake of simplicity let me consider \(a\) and \(c\) as the greater of the roots respectively. So, \(a−b=c−d\) Now, \(a−b=\sqrt{(a−b)^2}=\sqrt{(a+b)^2−4ab}=\sqrt{(−p)^2−4(q)}=\sqrt{p^2−4q}\) Similarly, \(c−d=\sqrt{q^2−4p}\) Now, a-b and c-d are equal. So, \(\sqrt{q^2−4p} = \sqrt{p^2−4q}\) (though it wouldn't have made a difference really) Now... . . .

OpenStudy (anonymous):

@shubhamsrg how (p+q)(p-q) = -4..................??

OpenStudy (shubhamsrg):

since p^2 - q^2 = 4(q-p) =>(p+q)(p-q) = -4(p-q)

OpenStudy (anonymous):

how p^2-q^2 = 4(q-p)?? tell me after- (p+2q)(p-2q) = (q+2p)(q-2p)

OpenStudy (anonymous):

@nitz help me with the question i had sent to you.

OpenStudy (anonymous):

ya i was solving it........

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@nitz you can that here.

OpenStudy (anonymous):

*solve

OpenStudy (anonymous):

shubham word me solve kiya hai .attach kar rahin hun see it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@nitz but x+1>0 or x>-1 also the ans. has (-1,0)U(1,infinity)

OpenStudy (anonymous):

now this was domain of function

OpenStudy (anonymous):

if we see the inequality (x+1/x-1)>2^0 (x+1/x-1)>1 x+1>x-1

OpenStudy (anonymous):

Have you removed the log here??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but how?? the log is only in numerator not in denominator.

OpenStudy (anonymous):

\[\log_{2}(x+1)\div x-1 > 0\]

OpenStudy (anonymous):

when did you tell that

OpenStudy (anonymous):

again i have to solveeee

OpenStudy (anonymous):

i've written it in the message sent to you.

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

no it seemed to be over entire part

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ya got it

OpenStudy (anonymous):

@nitz what happened?

OpenStudy (anonymous):

sorry light chalee gayee thee

OpenStudy (anonymous):

koi baat nhi. Its India!!

OpenStudy (anonymous):

case 1: when x-1>0 ie when x>1 then,log(base 2)(x+1)>0 x+1>2^0 x+1>1 x>0 that is when x>1 then x>0..................(A)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

case 2: when x-1<0 ie x<1 log(base 2)(x+1)<0....................sign is changed x+1<1 x<0 now when x<1 then x<0

OpenStudy (anonymous):

also here domain is x+1>0 ie x>-1

OpenStudy (anonymous):

log(base 2)(x+1)<0....................sign is changed ??

OpenStudy (anonymous):

when denominator is negative,sign gets reversed

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so combining all these conditions,you get your answer

OpenStudy (anonymous):

the commom interval is (-1,0)U(1,inf)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

answer aa gaya

OpenStudy (anonymous):

trying!!

OpenStudy (anonymous):

ok tell me where you dont get??

OpenStudy (anonymous):

ok my combined answer is not coming.

OpenStudy (anonymous):

@nitz i am getting x =(-1,0)U(0,1)U(1,inf.)

OpenStudy (anonymous):

see in first case the common interval will be \[(1,\infty) \]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok got it. I had taken the intersection of all 5 conditions.

OpenStudy (anonymous):

in second case the common interval was to be \[(-\infty,0) \] but since x>-1 so common interval is (-1,0)

OpenStudy (anonymous):

so (-1,0)U(1,inf)

OpenStudy (anonymous):

interested in quadratics??

OpenStudy (anonymous):

i got some problems there too.